If simple Keynesian model with no government,taxes or foreign sector
C=100+0,9y
I=50
1. What is the value of the multiplier in this economy?
2. Suppose the level of output that created full employment is 1800.Using the multiplier determine the is the level of investment spending that would create full employment in this economy.
1
Expert's answer
2019-08-13T08:57:24-0400
  1. The value of the marginal propensity to consume is: MPC = 0.9., then the value of the multiplier in this economy is:

1/(1-0.9) = 10

2 / The equilibrium level of output is:

Y = C+ I = 100 + 0.9Y + 50,

0.1Y= 150,

(1800-1500)/(I-50) = 10

I - 50 = 30

I = 80 - the level of investment spending that would create full employment in this economy